Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Monday, November 16, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 16, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/16/98 (Monday)

Segment 1 (Emidio to ANF) Report for 16 November, 1998. A crew prepared to push drying pigs from Copco road, EMP 33.34 to the "Edison Skip" on Segment 2. An excavator dug a pit at the Cuddy creek tributary culverts on Peace Valley road, EMP 25.51. A bell hole was excavated at ~EMP 21.50, to expose the end of previously buried pipe. Approximately 50 feet of trench was opened from this bell hole to the south. Trench was backfilled at EMP 18.5. An excavator continued to remove rock and open trench near EMP 18.14. On Grapevine Road, Excavation of the trench has progressed to ~EMP 15.59. Bell holes have been excavated to ~EMP 15.84. The coating crew finished near EMP 16.20. The Pipe bending operation moved to EMP 16.00, where a section of ROW was cleared of vegetation to accommodate the bending machine. A water truck was used to suppress fugitive dust along Grapevine road in the afternoon. Backfill, recontouring took place along Aqueduct Access road B. Bore equipment was delivered to the road B crossing at EMP 6.0. Trenching continued along the north shoulder of Hwy 166, in the area of EMP 1. West of the trenching, backfill of ditch continued. Welding and electrical crews continue to work at both Emidio and Grapevine Stations.

Segment 2 (ANF to Sylmar) Report for 11/15/98: 2A (Copco Rd to Templin Hwy): The drying 'pig' device that was pushed from the Copco Rd testing cutout towards Segment 2C on 11/13/98 had stalled because of generator difficulties. This cleaning run was started up again today. The drying 'pig' was scheduled to arrive at the Marple Cyn cutout, EMP 55.46, during the late evening. 2C (Templin Hwy to Castaic): Recontouring continued at the Templin Hwy cutout. Soil was brought in to this site from the EMP 55.96 spoils overrun to aid in the recontouring. Steel plates used to stabilize the ANF access road bed south of Whitaker Reduction Station near EMP 51.90 during construction were removed. Erosion protection work took place at gullies crossing the ROW between EMP 53.60 and 54.41. Spoils were removed from the base of the overrun at EMP 55.96 in preparation for construction of a block wall to protect the access road below. A crew manned the exit cutout site at EMP 55.46 for the drying 'pig' run from Segment 2A . Recontouring moved north towards the EMP 55.46 cutout to 55.70. General clean-up work continued throughout Segment 2C. 2B (Castaic to Sylmar): Crews continued pulling F/O cable from the Castaic yard on Castaic Road to Lake Hughes Parkway. At Rye Cyn, electrical work continued on the vault. At the directional drill at Feedmill Rd, drilling of the F/O bore reached 500 feet. Crews worked to control a frac-out of drilling mud that occurred during the weekend. Mud was pumped out, and a sandbag berm was constructed around the frac-out for containment. At Calgrove, tie-in was completed at the south end of the bore. Trenching occurred south of Weldon Cyn at EMP 72.45. At Sunshine Cyn, asphalt was completed through the intersection, although some of this section may have to be redone.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 16 November 1998: 3A: Backfill, padding, and fiber optic crews worked on closing up the ditch at the San Fernando Road Bridge crossing. Welding, x-ray, and coating crews worked at Hollywood Way Bridge crossing as well as on the bore at Van Nuys Boulevard. A crew also dug out part of the collapsed ditch caused by a car wrecking into the ditch over the weekend. The hangers were all placed on Tujunga Wash Bridge crossing. Electrical work continued on the valve north of Tujunga Wash. Holes were drilled on the bridge at the Pacoima Wash crossing. 3B: No work occurred. 3C: Work was done on digging out the exit hole for the Hubbard bore. General clean up was done along the right of way. The tie in from the road back to the railroad right of way neared completion today, MP 74.29. Crews began digging out the small skip between the railroad bore and the aqueduct crossing, to MP 73.99.

Segment 4B(El Segundo to Lynwood) 11-16-98: The crew at El Segundo Station ESEMP 11.00 continued to work on the 90 degree tie-in. The crew at Western ESEMP 5.56 was backfilling. The crew inside Lynwood station were working on the electrical wires and the back wall. The crew at College St. was repairing the fiberoptic conduit.

Segment 5 (Lynwood to Long Beach) Report for 16 November 1998: Directional drill at Highway 91/Compton Creek continued. Cased bore and pothole at Wilmington refinery progressed.

Inspection Summary

In the past 24 hours, 5 inspection event(s) were reported for 5 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page